The ratio of the measures of two complimentary angles is 3:7. What is the measure of the larger angle?
Answer by mananth(16949)

angle = x
complementary angle = 90-x
x/(90-x)=3/7
cross multiply
7x=3(90-x)
7x=210-3x
10x=210
x=21 deg
complementary =69 deg
m.ananth@hotmail.ca